OpenStudy (anonymous):

Identify the first step using the elimination method to solve the following system of equations: 10x - 7y = -18 15x + 3y = 27

OpenStudy (anonymous):

A. 7(10x - 7y = -18) 3(15x + 3y = 27) B. 15(10x - 7y = -18) 10(15x + 3y = 27) C. -3(10x - 7y = -18) 2(15x + 3y = 27) D. 27(10x - 7y = -18) 18(15x + 3y = 27)